Steam is the working fluid in a simple, ideal Rankine cycle. Saturated vapor enters the turbine at 8 MPa and saturated liquid exits the condenser at a pressure of 8 kPa. The net power output of the cycle is 100 MW. Determine for the cycle:

i. Thermal efficiency

ii. Back work ratio

iii. Mass flow-rate of the steam in kg/h

iv. Rate of heat transfer to the working fluid as it passes through the boiler in MW

v. Rate of heat transfer from the working fluid as is passes through the condenser in MW

vi. The mass flow rate of the condenser cooling water in kg/h if the cooling water enters the condenser at 15
Answer #1
Concepts and reason

Ideal Rankine cycle:

An ideal Rankine cycle consists of the following four processes:

1-2: Isentropic expansion in turbine

2-3: Isothermal heat rejection in condenser

3-4: Isentropic compression in pump

4-1: Isothermal heat addition in boiler

The T-s diagram for an ideal Rankine cycle is as follows:

T
/8000 kPa
(1)
4
8 kPa
2
3

Fundamentals

Write the equation for the thermal efficiency of the cycle is:

η=W˙tW˙pQ˙in\eta = \frac{{{{\dot W}_t} - {{\dot W}_p}}}{{{{\dot Q}_{in}}}}

Here, the work done by the turbine is W˙t{\dot W_t} , the work done on the pump is W˙p{\dot W_p} , and the heat input to the boiler is Q˙in{\dot Q_{in}} .

Write the equation for the back work ratio:

bwr=W˙pW˙t{\rm{bwr}} = \frac{{{{\dot W}_p}}}{{{{\dot W}_t}}}

The mass flow rate of the steam is as follows::

m˙=W˙cycleW˙tW˙p\dot m = \frac{{{{\dot W}_{{\rm{cycle}}}}}}{{{{\dot W}_t} - {{\dot W}_p}}}

The rate of heat transfer as the fluid passes through boiler is as follows:

Q˙in=m˙(h1h4){\dot Q_{in}} = \dot m\left( {{h_1} - {h_4}} \right)

The rate of heat transfer as the fluid passes through condenser is as follows:

Q˙out=m˙(h2h3){\dot Q_{out}} = \dot m\left( {{h_2} - {h_3}} \right)

State 1:

Obtain the properties corresponding to p1=8MPa{p_1} = 8{\rm{ MPa}} and x1=1{x_1} = 1 (saturated vapour) from the Table, "Saturated water- Pressure table."

Specific enthalpy at state 1, h1=hg=2758.0kJ/kg{h_1} = {h_g} = 2758.0{\rm{ kJ/kg}}

Specific entropy at state 1, s1=sg=5.7432kJ/kgK{s_1} = {s_g} = 5.7432{\rm{ kJ/kg}} \cdot {\rm{K}}

State 2:

Since, the process 1-2 is isentropic,

s2=s1=5.7432kJ/kgK\begin{array}{c}\\{s_2} = {s_1}\\\\ = 5.7432{\rm{ kJ/kg}} \cdot {\rm{K}}\\\end{array}

Obtain the properties corresponding to p2=8kPa{p_2} = 8{\rm{ kPa}} and s2=5.7432kJ/kgK{s_2} = 5.7432{\rm{ kJ/kg}} \cdot {\rm{K}} (saturated vapour) from the Table, "Saturated water- Pressure table."

Specific entropy of saturated liquid at state 2, sf=0.5926kJ/kgK{s_f} = 0.5926{\rm{ kJ/kg}} \cdot {\rm{K}}

Specific entropy of liquid-vapour phase at state 2, sfg=7.6361kJ/kgK{s_{fg}} = 7.6361{\rm{ kJ/kg}} \cdot {\rm{K}}

Specific enthalpy of saturated liquid at state 2, hf=173.88kJ/kg{h_f} = 173.88{\rm{ kJ/kg}}

Specific enthalpy liquid-vapour phase at state 2, hfg=2403.1kJ/kg{h_{fg}} = 2403.1{\rm{ kJ/kg}}

Calculate the quality of steam at state 2:

s2=sf+x2sfg{s_2} = {s_f} + {x_2}{s_{fg}}

Substitute 5.7432kJ/kgK5.7432{\rm{ kJ/kg}} \cdot {\rm{K}} s2{s_2} , 0.5926kJ/kgK0.5926{\rm{ kJ/kg}} \cdot {\rm{K}} for sf{s_f} , and 7.6361kJ/kgK7.6361{\rm{ kJ/kg}} \cdot {\rm{K}} for sfg{s_{fg}} .

5.7432=0.5926+x2(7.6361)x2=5.74320.59267.6361x2=0.6745\begin{array}{l}\\5.7432 = 0.5926 + {x_2}\left( {7.6361} \right)\\\\{x_2} = \frac{{5.7432 - 0.5926}}{{7.6361}}\\\\{x_2} = 0.6745\\\end{array}

Calculate the specific enthalpy at state 2:

h2=hf+x2hfg{h_2} = {h_f} + {x_2}{h_{fg}}

Substitute 173.88kJ/kg173.88{\rm{ kJ/kg}} for hf{h_f} , 0.67450.6745 for x2{x_2} , and 2403.1kJ/kg2403.1{\rm{ kJ/kg}} for hfg{h_{fg}} .

h2=173.88+(0.6745)(2403.1)=1794.8kJ/kg\begin{array}{c}\\{h_2} = 173.88 + \left( {0.6745} \right)\left( {2403.1} \right)\\\\ = 1794.8{\rm{ kJ/kg}}\\\end{array}

State 3:

Obtain the properties corresponding to p3=8kPa{p_3} = 8{\rm{ kPa}} and x1=0{x_1} = 0 (saturated liquid) from the Table, "Saturated water- Pressure table."

Specific enthalpy at state 3, h3=hf=173.88kJ/kg{h_3} = {h_f} = 173.88{\rm{ kJ/kg}}

Specific volume at state 3, v3=vf=1.0085×103m3/kg{v_3} = {v_f} = 1.0085 \times {10^{ - 3}}{\rm{ }}{{\rm{m}}^3}{\rm{/kg}}

State 4:

Write the relation for the pump work:

W˙pm˙=h4h3h4=W˙pm˙+h3h4=v3(p4p3)+h3\begin{array}{l}\\\frac{{{{\dot W}_p}}}{{\dot m}} = {h_4} - {h_3}\\\\{h_4} = \frac{{{{\dot W}_p}}}{{\dot m}} + {h_3}\\\\{h_4} = {v_3}\left( {{p_4} - {p_3}} \right) + {h_3}\\\end{array}

Substitute 173.88kJ/kg173.88{\rm{ kJ/kg}} for h3{h_3} , 1.0085×103m3/kg1.0085 \times {10^{ - 3}}{\rm{ }}{{\rm{m}}^3}{\rm{/kg}} for v3{v_3} , 8000kPa8000{\rm{ kPa}} for p4{p_4} , and 8kPa8{\rm{ kPa}} for p3{p_3} .

h4=1.0085×103(80008)+173.88=181.94kJ/kg\begin{array}{c}\\{h_4} = 1.0085 \times {10^{ - 3}}\left( {8000 - 8} \right) + 173.88\\\\ = 181.94{\rm{ kJ/kg}}\\\end{array}

Calculate the thermal efficiency of the cycle.

η=W˙tW˙pQ˙in=m˙(h1h2)m˙(h4h3)m˙(h1h4)=(h1h2)(h4h3)(h1h4)\begin{array}{c}\\\eta = \frac{{{{\dot W}_t} - {{\dot W}_p}}}{{{{\dot Q}_{in}}}}\\\\ = \frac{{\dot m\left( {{h_1} - {h_2}} \right) - \dot m\left( {{h_4} - {h_3}} \right)}}{{\dot m\left( {{h_1} - {h_4}} \right)}}\\\\ = \frac{{\left( {{h_1} - {h_2}} \right) - \left( {{h_4} - {h_3}} \right)}}{{\left( {{h_1} - {h_4}} \right)}}\\\end{array}

Substitute 2758.0kJ/kg2758.0{\rm{ kJ/kg}} for h1{h_1} , 1794.8kJ/kg1794.8{\rm{ kJ/kg}} for h2{h_2} , 173.88kJ/kg173.88{\rm{ kJ/kg}} for h3{h_3} , and 181.94kJ/kg181.94{\rm{ kJ/kg}} for h4{h_4} .

η=(27581794.8)(181.94173.88)(2758.0181.94)=0.371=37.1%\begin{array}{c}\\\eta = \frac{{\left( {2758 - 1794.8} \right) - \left( {181.94 - 173.88} \right)}}{{\left( {2758.0 - 181.94} \right)}}\\\\ = 0.371\\\\ = 37.1\% \\\end{array}

Calculate the back work ratio.

bwr=W˙pW˙t=h4h3h1h2\begin{array}{c}\\{\rm{bwr}} = \frac{{{{\dot W}_p}}}{{{{\dot W}_t}}}\\\\ = \frac{{{h_4} - {h_3}}}{{{h_1} - {h_2}}}\\\end{array}

Substitute 2758.0kJ/kg2758.0{\rm{ kJ/kg}} for h1{h_1} , 1794.8kJ/kg1794.8{\rm{ kJ/kg}} for h2{h_2} , 173.88kJ/kg173.88{\rm{ kJ/kg}} for h3{h_3} , and 181.94kJ/kg181.94{\rm{ kJ/kg}} for h4{h_4} .

bwr=181.94173.88227581794.8=8.4×103=84%\begin{array}{c}\\{\rm{bwr}} = \frac{{181.94 - 173.882}}{{2758 - 1794.8}}\\\\ = 8.4 \times {10^{ - 3}}\\\\ = 84\% \\\end{array}

Calculate the mass flow rate of the steam.

m˙=W˙cycleW˙tW˙p=W˙cycle(h1h2)(h4h3)\begin{array}{c}\\\dot m = \frac{{{{\dot W}_{{\rm{cycle}}}}}}{{{{\dot W}_t} - {{\dot W}_p}}}\\\\ = \frac{{{{\dot W}_{{\rm{cycle}}}}}}{{\left( {{h_1} - {h_2}} \right) - \left( {{h_4} - {h_3}} \right)}}\\\end{array}

Here, the work done during the cycle is W˙cycle{\dot W_{{\rm{cycle}}}} .

Substitute 110MW110{\rm{ MW}} for W˙cycle{\dot W_{{\rm{cycle}}}} , 2758.0kJ/kg2758.0{\rm{ kJ/kg}} for h1{h_1} , 1794.8kJ/kg1794.8{\rm{ kJ/kg}} for h2{h_2} , 173.88kJ/kg173.88{\rm{ kJ/kg}} for h3{h_3} , and 181.94kJ/kg181.94{\rm{ kJ/kg}} for h4{h_4} .

m˙=W˙cycleW˙tW˙p=110MW×(1000kW1MW)×(1kJ/s1kW)×(3600s1h)(27581794.8)(181.94173.88)=3.77×105kg/h\begin{array}{c}\\\dot m = \frac{{{{\dot W}_{{\rm{cycle}}}}}}{{{{\dot W}_t} - {{\dot W}_p}}}\\\\ = \frac{{110{\rm{ MW}} \times \left( {\frac{{1000{\rm{ kW}}}}{{1{\rm{ MW}}}}} \right) \times \left( {\frac{{1{\rm{ kJ/s}}}}{{1{\rm{ kW}}}}} \right) \times \left( {\frac{{3600{\rm{ s}}}}{{1{\rm{ h}}}}} \right)}}{{\left( {2758 - 1794.8} \right) - \left( {181.94 - 173.88} \right)}}\\\\ = 3.77 \times {10^5}{\rm{ kg/h}}\\\end{array}

Calculate the rate of heat transfer to the fluid upon passing through the boiler.

Q˙in=m˙(h1h4){\dot Q_{in}} = \dot m\left( {{h_1} - {h_4}} \right)

Substitute 3.77×105kg/h3.77 \times {10^5}{\rm{ kg/h}} for m˙\dot m , 2758.0kJ/kg2758.0{\rm{ kJ/kg}} for h1{h_1} , and 181.94kJ/kg181.94{\rm{ kJ/kg}} for h4{h_4} .

Q˙in=3.77×105×(13600)kg/s×(2758.0181.94)kJkg=269.77×103kW=269.77MW\begin{array}{c}\\{{\dot Q}_{in}} = 3.77 \times {10^5} \times \left( {\frac{1}{{3600}}} \right){\rm{ kg/s}} \times \left( {2758.0 - 181.94} \right){\rm{ }}\frac{{{\rm{kJ}}}}{{{\rm{kg}}}}\\\\ = 269.77 \times {10^3}{\rm{ kW}}\\\\ = 269.77{\rm{ MW}}\\\end{array}

Calculate the rate of heat transfer to the fluid upon passing through the condenser.

Q˙out=m˙(h2h3){\dot Q_{out}} = \dot m\left( {{h_2} - {h_3}} \right)

Substitute 3.77×105kg/h3.77 \times {10^5}{\rm{ kg/h}} for m˙\dot m , 1794.8kJ/kg1794.8{\rm{ kJ/kg}} for h2{h_2} , and 173.88kJ/kg173.88{\rm{ kJ/kg}} for h3{h_3} .

Q˙out=3.77×105×(13600)kg/s×(1794.8173.88)kJkg=169.75×103kW=169.75MW\begin{array}{c}\\{{\dot Q}_{out}} = 3.77 \times {10^5} \times \left( {\frac{1}{{3600}}} \right){\rm{ kg/s}} \times \left( {1794.8 - 173.88} \right){\rm{ }}\frac{{{\rm{kJ}}}}{{{\rm{kg}}}}\\\\ = 169.75 \times {10^3}{\rm{ kW}}\\\\ = 169.75{\rm{ MW}}\\\end{array}

Obtain the properties corresponding to Tin=15C{T_{{\rm{in}}}} = 15^\circ {\rm{C}} and x1=0{x_1} = 0 (saturated liquid) from the Table, "Saturated water- temperature table."

Specific enthalpy of inlet cooling water, hc,in=hf=62.99kJ/kg{h_{c,in}} = {h_f} = 62.99{\rm{ kJ/kg}}

Obtain the properties corresponding to Tout=35C{T_{out}} = 35^\circ {\rm{C}} and x1=0{x_1} = 0 (saturated liquid) from the Table, "Saturated water- temperature table."

Specific enthalpy of outlet cooling water, hc,out=hf=146.68kJ/kg{h_{c,out}} = {h_f} = 146.68{\rm{ kJ/kg}}

Calculate the mass flow rate of the condenser cooling water by using steady flow energy equation:

0=Q˙cvW˙cv+m˙c[hc,inhc,out]+m˙(h2h3)m˙c[hc,inhc,out]+Q˙out=0m˙c=Q˙outhc,outhc,in\begin{array}{l}\\0 = {{\dot Q}_{cv}} - {{\dot W}_{cv}} + {{\dot m}_c}\left[ {{h_{c,in}} - {h_{c,out}}} \right] + \dot m\left( {{h_2} - {h_3}} \right)\\\\{{\dot m}_c}\left[ {{h_{c,in}} - {h_{c,out}}} \right] + {{\dot Q}_{out}} = 0\\\\{{\dot m}_c} = \frac{{{{\dot Q}_{out}}}}{{{h_{c,out}} - {h_{c,in}}}}\\\end{array}

Here, the mass flow rate of the condenser cooling water is m˙c{\dot m_c} .

Substitute 169.75MW169.75{\rm{ MW}} for Q˙out{\dot Q_{out}} , 62.99kJ/kg62.99{\rm{ kJ/kg}} for hc,in{h_{c,in}} , and 146.68kJ/kg146.68{\rm{ kJ/kg}} for hc,out{h_{c,out}} .

m˙c=169.75×103kW(146.6862.99)kJ/kg×(1kJ/s1kW)×(3600s1h)=7.3×106kg/h\begin{array}{c}\\{{\dot m}_c} = \frac{{169.75 \times {{10}^3}{\rm{ kW}}}}{{\left( {146.68 - 62.99} \right){\rm{ kJ/kg}}}} \times \left( {\frac{{1{\rm{ kJ/s}}}}{{1{\rm{ kW}}}}} \right) \times \left( {\frac{{3600{\rm{ s}}}}{{1{\rm{ h}}}}} \right)\\\\ = 7.3 \times {10^6}{\rm{ kg/h}}\\\end{array}

Ans: Part i

The thermal efficiency of the cycle is 37.1%{\bf{37}}{\bf{.1\% }} .

Part ii

The back work ratio is 84%{\bf{84\% }} .

Part iii

The mass flow rate of the steam is 3.77×105kg/h{\bf{3}}{\bf{.77 \times 1}}{{\bf{0}}^{\bf{5}}}{\bf{ kg/h}} .

Part iv

The rate of heat transfer to the fluid upon passing through the boiler is 269.77W{\bf{269}}{\bf{.77 W}} .

Part v

The rate of heat transfer to the fluid upon passing through the condenser is 169.75W{\bf{169}}{\bf{.75 W}} .

Part vi

The mass flow rate of the condenser cooling water is 7.3×106kg/h{\bf{7}}{\bf{.3 \times 1}}{{\bf{0}}^{\bf{6}}}{\bf{ kg/h}} .